# Cross Validation in Stochastic Analytic Continuation

### v2

@@ A.1 $A_{0}$ Scan
Figure 14 shows the results of the $A_{0}$ scans used to determine the optimal weight of the macroscopic $\delta$-peak for the two momenta considered, $q=\pi/2$, panel (a), and $q=3\pi/4$, panel (b). In this case, unrestricted sampling runs (i.e. $A_{0}=0$) were first preformed to determine the minimum $\chi^{2}$ values reached at the end of each anneal. This value is then used to determine the sampling temperature for the $A_{0}>0$ runs; we use a value of $\Theta$ that is slightly elevated from that corresponding to the criterion Eq. (11) for this unrestricted sampling run [20]. At this temperature, a clear minimum is present, but the $\chi^{2}$ value is still at an acceptable value. The optimal values of $A_{0}$, as determined by the results shown in Fig. 14, are $A_{0}=0.40$ and $0.35$ for $q=\pi/2$ and $q=3\pi/4$, respectively.
Figure 15: Variable amplitudes used to produce the $\delta$-function diagram in Fig. 1(d). For an exponent $p=-0.750$, the amplitudes increase as the edge is approached. The location of the equal amplitude cross-over point, $n_{0}$, is marked.

### v3

@@ Appendix B Optimal SAC Parameters for the Heisenberg Chain
When implementing the single-peak and edge SAC parameterizations, a fixed parameter must be chosen using a simple optimization procedure. In the case of the single-peak parameterization, this parameter is the weight of the macroscopic $\delta$-peak, $A_{0}$, and in the case of the edge parameterization, this parameter is the power $p$ with which the edge asymptotically diverges with. Scans over $A_{0}$ and $p$ are performed before cross validation to determine their optimal values.
In either case, the SAC annealing routine in run multiple times with different fixed values of $A_{0}$ or $p$, where typically the values are selected from a sensibly defined grid (not too dense or too sparse). After these scans are performed, the $\chi^{2}$ values as a function of either $A_{0}$ or $p$ are compared at a fixed value of the sampling temperature and a minimum is located. At very low values of $\Theta$, where the sampling is dominated by noise in the QMC-generated data, the $\chi^{2}$ minimum will be very shallow, if present at all, so instead an elevated temperature is used. This temperature can be chosen based on the minimum $\chi^{2}$ reach in an unrestricted sampling SAC run on the same set of data, as detailed in Ref. 20, or simply set to a value high enough for a clear minimum to emerge (but low enough so that the $\chi^{2}$ value at the minimum is still acceptable). The exact location of the minimum depends on the sampling temperature to some extent, but in most cases this dependance is very minor. Furthermore, any value of $A_{0}$ or $p$ chosen around the $\chi^{2}$ minimum will produce nearly identical spectra.
It is crucial to have as highest possible data quality when determining the optimal values of $A_{0}$ and $p$. Therefore, we use the full set of QMC-generated data for the Heisenberg chain when performing the parameter scans. In this case, the complete $G(\tau)$ data set has an error level $\sim~{}3~{}\times~{}10^{-6}$, nearly ten times smaller than the error level of the validation data sets.
B.1 $A_{0}$ Scan
Figure 15 shows the results of the $A_{0}$ scans used to determine the optimal weight of the macroscopic $\delta$-peak for the two momenta considered, $q=\pi/2$, panel (a), and $q=3\pi/4$, panel (b). In this case, unrestricted sampling runs (i.e. $A_{0}=0$) were first performed to determine the minimum $\chi^{2}$ values reached at the end of each anneal. This value is then used to determine the sampling temperature for the $A_{0}>0$ runs; we use a value of $\Theta$ that is slightly elevated from that corresponding to the criterion Eq. (11) for this unrestricted sampling run [20]. At this temperature, a clear minimum is present, but the $\chi^{2}$ value is still at an acceptable value. The optimal values of $A_{0}$, as determined by the results shown in Fig. 15, are $A_{0}=0.40$ and $0.35$ for $q=\pi/2$ and $q=3\pi/4$, respectively.
Figure 16: Variable amplitudes used to produce the $\delta$-function diagram in Fig. 1(d). For an exponent $p=-0.75$, the amplitudes increase as the edge is approached. The location of the equal amplitude cross-over point, $n_{0}$, is marked.